# Solving the Granularity Mismatch: Hierarchical Preference Learning for Long-Horizon LLM Agents

### v1

@@ Abstract
Large Language Models (LLMs) as autonomous agents are increasingly tasked with solving complex, long-horizon problems. Aligning these agents via preference-based offline methods like Direct Preference Optimization (DPO) is a promising direction, yet it faces a critical granularity mismatch. Trajectory-level DPO provides a signal that is too coarse for precise credit assignment, while step-level DPO is often too myopic to capture the value of multi-step behaviors. To resolve this challenge, we introduce Hierarchical Preference Learning (HPL), a hierarchical framework that optimizes LLM agents by leveraging preference signals at multiple, synergistic granularities. While HPL incorporates trajectory- and step-level DPO for global and local policy stability, its core innovation lies in group-level preference optimization guided by a dual-layer curriculum. HPL first decomposes expert trajectories into semantically coherent action groups and then generates contrasting suboptimal groups to enable preference learning at a fine-grained, sub-task level. Then, instead of treating all preference pairs equally, HPL introduces a curriculum scheduler that organizes the learning process from simple to complex. This curriculum is structured along two axes: the group length, representing sub-task complexity, and the sample difficulty, defined by the reward gap between preferred and dispreferred action groups. Experiments on three challenging agent benchmarks show that HPL outperforms existing state-of-the-art methods. Our analyses demonstrate that the hierarchical DPO loss effectively integrates preference signals across multiple granularities, while the dual-layer curriculum is crucial for enabling the agent to solve a wide range of tasks, from simple behaviors to complex multi-step sequences.
@@ 1 Introduction
To equip agents for such tasks, Reinforcement Learning (RL) has become a crucial recipe for post-training LLMs (Liu et al., 2024). Online RL methods like PPO (Sutton et al., 1998; Schulman et al., 2017) often entail substantial computational costs, high sample inefficiency, and risky, inefficient exploration in vast action spaces. These challenges have motivated a growing interest in offline RL paradigms that learn from static datasets. Direct Preference Optimization (DPO) (Rafailov et al., 2023) directly aligns agent policies using preference pairs (e.g., expert vs. suboptimal behaviors) without requiring costly environment interaction or an explicitly trained reward model.
However, applying DPO to long-horizon agent tasks reveals a fundamental challenge we term the granularity mismatch. On one hand, trajectory-level DPO, such as ETO (Song et al., 2024), compares entire trajectories. This approach provides a learning signal that is often too coarse and sparse; it fails to pinpoint which specific sequence of actions within a long history was responsible for the outcome. On the other hand, step-level DPO, employed by methods such as IPR (Xiong et al., 2024), operates at the finest granularity by comparing the consequences of single actions. However, this fine-grained focus presents a challenge: it can struggle to capture the synergistic value of actions that are only meaningful as part of a longer sequence. For instance, the sub-task of “retrieving an apple from the fridge” is composed of a chain of actions—navigating, opening, and taking—whose collective value cannot be captured by rewarding any single action in isolation.
@@
To resolve this dilemma, we introduce Hierarchical Preference Learning (HPL), a hierarchical framework that optimizes LLM agents by leveraging preference signals at multiple, synergistic granularities. HPL first addresses the granularity mismatch by incorporating DPO losses at the trajectory, action, and the action-group levels. These groups, which often correspond to functional sub-tasks, provide an ideal intermediate unit for credit assignment. Beyond merely combining these losses, the core innovation of HPL is a dual-layer curriculum learning strategy that guides the training process. This curriculum systematically organizes the learning path from simple to complex along two orthogonal axes: sub-task complexity, defined by the length of an action group, and sample difficulty, measured by the reward gap between preferred and dispreferred behaviors. By first mastering simple, easily distinguishable sub-tasks, the agent builds a foundation before progressing to more complex challenges.
@@
Figure 1: Conceptual comparison of different DPO granularities. While (a) trajectory-level DPO provides a coarse signal and (b) step-level DPO provides a myopic one, (c) our proposed Group-level DPO learns from semantically coherent action groups, which provides a structured and meaningful signal, enabling the agent to reason at the sub-task level.
@@
  • We identify and address the granularity mismatch problem in offline agent alignment by proposing a novel hierarchical framework that integrates preference signals at three distinct levels: the coarse trajectory-level, the fine-grained step-level, and a intermediate action-group level.
@@
  • We demonstrate through extensive experiments on three diverse and challenging long-horizon benchmarks that HPL significantly outperforms existing state-of-the-art methods, establishing a more effective and principled paradigm for offline LLM agent alignment.
@@ 2 Related Work
LLM-based Agents. The remarkable reasoning and instruction-following capabilities of modern LLMs have enabled their use as autonomous agents capable of tackling complex, interactive tasks (Li et al., 2024; Gou et al., 2025). Initial approaches primarily leveraged the in-context learning abilities of LLMs through prompts like ReAct (Yao et al., 2023) and Reflexion (Shinn et al., 2023) to elicit multi-step reasoning and action generation. To enhance the performance of open-source models beyond zero-shot prompting, a subsequent line of work has focused on fine-tuning agents on collected trajectory data (Chen et al., 2023). These methods range from standard Supervised Fine-Tuning (SFT) on expert demonstrations (Zeng et al., 2023; Chen et al., 2024) to more advanced techniques that learn from preference data, such as contrasting successful and failed trajectories to optimize for final outcomes (Song et al., 2024). While effective, these fine-tuning paradigms often treat entire trajectories as monolithic data points (Wang et al., 2024b), overlooking the fine-grained procedural knowledge embedded within the interaction trajectory.
Process Supervision. To address the limitations of outcome-based rewards, particularly the challenge of credit assignment in long-horizon tasks, a growing body of research has explored process supervision (Luo et al., 2024; Xiong et al., 2025). The core idea is to provide agents with more granular feedback at intermediate steps of a task. Early efforts in this area often relied on costly human annotations to label the correctness of each step (Lightman et al., 2023). To automate this, recent methods have proposed various techniques to estimate step-level rewards, such as using Monte Carlo rollouts to predict the future outcome from an intermediate state (Xiong et al., 2024) or training a separate reward model to predict the value of each action (Choudhury, 2025; Wang et al., 2025). These approaches typically use the estimated step-level rewards to guide the agent via reinforcement learning (Feng et al., 2025; Zhang et al., 2025) or DPO at the single-action level.
@@ 3.2.2 Group-Level Data Generation via Action Group Segmentation
To bridge the gap between coarse trajectories and myopic single actions, we introduce the core concept of an action group. These groups serve as an intermediate unit of reasoning, ideally corresponding to semantically coherent sub-tasks for more effective credit assignment. The generation of group-level data involves two key steps: segmenting trajectories into groups and then estimating a quantitative reward for each group.
@@ 3.4 Multi-Granularity Preference Optimization
To better understand the effectiveness of the group-level objective, we briefly analyze its bias-variance properties in the following proposition, with a full derivation provided in Appendix D.
@@
Thus, the final training objective of our HPL combines these losses:
$\mathcal{L}_{\text{HPL}}^{(s)}=\mathcal{L}_{\text{BC}}+\mathcal{L}_{\text{traj-DPO}}+\mathcal{L}_{\text{step-DPO}}+\mathcal{L}_{\text{group-DPO}}^{(s)},$ (8)
@@ 4.1 Experimental Setup
We evaluate our proposed method, HPL, on three diverse and challenging long-horizon agent benchmarks: ALFWorld (Shridhar et al., 2021), WebShop (Yao et al., 2022), and InterCode-SQL (Yang et al., 2023). For all experiments, we use Qwen2.5-1.5B-Instruct and Qwen2.5-7B-Instruct as the backbone language models. HPL is compared against a suite of strong baselines, including SFT, RFT (Yuan et al., 2023), ETO (Song et al., 2024), and IPR (Xiong et al., 2024). All methods are initialized from an SFT model trained on expert trajectories generated by a GPT-4o teacher model. For Fixed-N and Fixed-K strategy, we set $N=3$ and $K=3$. For Uncertainty strategy, the entropy threshold $\epsilon$ is set to the 80th percentile of all action entropies (i.e., the threshold for the top 20% highest values) computed across the training dataset. Detailed descriptions of the environments, baseline implementations, and all hyperparameters are deferred to Appendix C.
@@ 4.2 Main Results (RQ1, RQ2)
Table 1: Performance comparison of HPL and strong baselines across three agent benchmarks. All methods are evaluated using Qwen2.5-1.5B-Instruct and Qwen2.5-7B-Instruct as base models. The best and second-best results are highlighted in bold and with an underline, respectively.
@@
   SFT 62.14 58.21 55.99 25.50 56.85 55.00 52.28
   RFT (Yuan et al., 2023) 61.43 60.45 57.88 29.00 57.99 56.50 53.88
   ETO (Song et al., 2024) 62.14 64.18 56.82 28.50 58.94 57.50 54.68
   IPR (Xiong et al., 2024) 64.29 64.93 58.83 29.00 59.74 58.00 55.64
@@
   SFT 67.14 76.12 60.53 32.00 66.86 65.00 61.28
   RFT (Yuan et al., 2023) 72.86 69.40 61.76 33.50 68.43 66.50 62.08
   ETO (Song et al., 2024) 70.00 76.87 61.57 33.50 68.63 67.00 62.93
   IPR (Xiong et al., 2024) 72.86 77.61 61.90 34.00 69.19 67.50 63.84
@@
As shown in Table 1, our HPL framework significantly outperforms all baseline methods across both model scales, providing a strong affirmative answer to our first research question. For the Qwen2.5-7B-Instruct model, our best-performing variant, HPL (Semantic), achieves an average score of 67.81, surpassing the strongest single-granularity baselines, ETO and IPR, by substantial margins of 4.88 and 3.97 points, respectively. This trend of superiority holds true across all three diverse and challenging benchmarks. The advantage of HPL’s hierarchical approach is particularly pronounced in tasks requiring complex, long-horizon generalization. For instance, on the ALFWorld benchmark’s unseen scenarios, HPL (Semantic) achieves a success rate of 86.57%, a remarkable improvement of nearly 9 points over the state-of-the-art IPR (77.61%). These results strongly indicate that by integrating preference signals at the trajectory, action, and group levels, HPL effectively resolves the granularity mismatch problem and learns a more robust and generalizable policy.
Our experiments also reveal the critical impact of the action group segmentation strategy, directly addressing our second research question. While all HPL variants consistently outperform the baselines, the adaptive, content-aware segmentation methods generally yield better performance than fixed heuristic approaches. Notably, HPL (Semantic), which leverages a powerful language model to partition trajectories into semantically coherent sub-tasks, consistently emerges as the top-performing variant for both the 1.5B and 7B models. For the 7B model, it outperforms the next-best variant, HPL (Fixed-N), by over 1 point in the average score. This suggests that the quality of the action groups is paramount; providing the DPO loss with more meaningful, human-aligned sub-tasks as units of comparison leads to a more effective and powerful learning signal. Even so, the strong performance of the simpler Uncertainty and Fixed-N strategies demonstrates the inherent value of incorporating an intermediate granularity, validating the core design of our HPL framework.
@@ 4.3 Analysis of the Curriculum Learning Mechanism (RQ3)
Ablation of Curriculum Components. To quantitatively assess the importance of our dual-layer curriculum, we conduct an ablation study by removing each layer individually, with results presented in Table 2. The primary finding is that the full HPL model, equipped with both curriculum layers, consistently outperforms all ablated variants. Removing the curriculum entirely (HPL Static) results in the most significant performance degradation across both model scales, confirming that employing a curriculum is crucial for effective learning. Furthermore, the results indicate that both the length and difficulty-based curricula contribute positively to the final performance, with their individual removal leading to noticeable performance drops. This demonstrates the synergistic benefit of our dual-layer design, where organizing the learning process first by task complexity (length) and then by solution quality (difficulty) provides a more effective path to mastering complex agent behaviors.

### v2

@@ Abstract
Large Language Models (LLMs) as autonomous agents are increasingly tasked with solving complex, long-horizon problems. Aligning these agents via preference-based methods like Direct Preference Optimization (DPO) is a promising direction, yet it faces a critical granularity mismatch. Trajectory-level DPO provides stable signals but blur where credit should be assigned within long trajectories, whereas step-level DPO offers fine-grained supervision but can be statistically noisy and data-inefficient when Monte Carlo rollouts are limited, and can be hard to fully exploit multi-step structured behaviors that only reveal their effect over several actions. To balance this trade-off, we introduce Hierarchical Preference Learning (HPL), a hierarchical framework that optimizes LLM agents by leveraging preference signals at multiple, complementary granularities. While HPL incorporates trajectory- and step-level DPO for global and local policy stability, its core innovation lies in group-level preference optimization guided by a dual-layer curriculum. HPL first decomposes expert trajectories into semantically coherent action groups and then generates contrasting suboptimal groups to enable preference learning at a fine-grained, sub-task level. Then, instead of treating all preference pairs equally, HPL introduces a curriculum scheduler that organizes the learning process from simple to complex. This curriculum is structured along two axes: the group length, representing sub-task complexity, and the sample difficulty, defined by the reward gap between preferred and dispreferred action groups. Experiments on three challenging agent benchmarks show that HPL outperforms existing state-of-the-art methods. Our analyses demonstrate that the hierarchical DPO loss effectively integrates preference signals across multiple granularities, while the dual-layer curriculum is crucial for enabling the agent to solve a wide range of tasks, from simple behaviors to complex multi-step sequences.
@@ 1 Introduction
To equip agents for such tasks, Reinforcement Learning (RL) has become a crucial recipe for post-training LLMs (Liu et al., 2024). Online RL methods like PPO (Sutton et al., 1998; Schulman et al., 2017) often entail substantial computational costs, high sample inefficiency, and risky, inefficient exploration in vast action spaces. These challenges have motivated approaches that reduce reliance on online interaction by learning from static datasets collected in advance. Direct Preference Optimization (DPO) (Rafailov et al., 2023) directly aligns agent policies using preference pairs (e.g., expert vs. suboptimal behaviors) without requiring costly environment interaction or an explicitly trained reward model.
However, applying DPO to long-horizon agent tasks reveals a fundamental challenge we term the granularity mismatch. On one hand, trajectory-level DPO, such as ETO (Song et al., 2024), compares entire trajectories and yields stable, low-variance feedback aligned with final outcomes, but it provides limited resolution for credit assignment, which is hard to tell which segment of a long interaction actually determined success or failure. On the other hand, step-level DPO, employed by methods such as IPR (Xiong et al., 2024), attributes preferences to individual decisions by estimating the expected return from each decision point via Monte Carlo rollouts. However, this fine-grained focus poses practical challenges in the finite-data, limited-rollout regime. Supervision becomes fragmented across many decision points, so each step is updated from only a few noisy rollouts, and it can be hard to fully exploit multi-step structured behaviors whose contribution to success only becomes apparent when several actions are considered jointly. For instance, the sub-task of “retrieving an apple from the fridge” is composed of a chain of actions—navigating, opening, and taking—whose collective value cannot be captured by rewarding any single action in isolation.
@@
To resolve this dilemma, we introduce Hierarchical Preference Learning (HPL), a hierarchical framework that optimizes LLM agents by leveraging preference signals at multiple, synergistic granularities. HPL first addresses the granularity mismatch by incorporating DPO losses at the trajectory, action, and the action-group levels. The group-level view provides both a structural prior by focusing supervision on sub-trajectories that are more likely to encode reusable skills and a statistical benefit by aggregating the contribution of multiple actions into one decision unit, which reduces variance relative to per-step Monte Carlo estimates under a fixed rollout budget. Beyond merely combining these losses, the core innovation of HPL is a dual-layer curriculum learning strategy that guides the training process. This curriculum systematically organizes the learning path from simple to complex along two orthogonal axes: sub-task complexity, defined by the length of an action group, and sample difficulty, measured by the reward gap between preferred and dispreferred behaviors. By first mastering simple, easily distinguishable sub-tasks, the agent builds a foundation before progressing to more complex challenges.
@@
Figure 1: Conceptual comparison of different DPO granularities. While (a) trajectory-level DPO provides a coarse but stable signal and (b) step-level DPO offers focused but potentially noisy supervision, (c) our proposed Group-level DPO learns from semantically coherent action groups, which provides a structured signal, enabling the agent to reason at the sub-task level.
@@
  • We identify and address the granularity mismatch problem in preference-based agent alignment by proposing a novel hierarchical framework that integrates preference signals at three distinct levels: the coarse trajectory-level, the fine-grained step-level, and a intermediate action-group level.
@@
  • We demonstrate through extensive experiments on three diverse and challenging long-horizon benchmarks that HPL significantly outperforms existing state-of-the-art methods, establishing a more effective and principled paradigm for preference-based training of LLM agents from fixed datasets after a one-shot exploration phase.
@@ 2 Related Work
LLM-based Agents. The remarkable reasoning and instruction-following capabilities of modern LLMs have enabled their use as autonomous agents capable of tackling a diverse array of complex, interactive tasks (Li et al., 2024; Gou et al., 2025; Tang et al., 2025; Wang et al., 2025b). Initial approaches primarily leveraged the in-context learning abilities of LLMs through prompts like ReAct (Yao et al., 2023) and Reflexion (Shinn et al., 2023) to elicit multi-step reasoning and action generation. This focus on inference-time computation has recently evolved into the broader paradigm of test-time scaling (TTS), where allocating additional compute during generation (Zhang et al., 2025b). However, to enhance the performance of open-source models beyond zero-shot prompting and and test-time strategies, a subsequent line of work has focused on fine-tuning agents on collected trajectory data (Chen et al., 2023). These methods range from standard Supervised Fine-Tuning (SFT) on expert demonstrations (Zeng et al., 2023; Chen et al., 2024) to more advanced techniques that learn from preference data, such as contrasting successful and failed trajectories to optimize for final outcomes (Song et al., 2024). While effective, these fine-tuning paradigms often treat entire trajectories as monolithic data points (Wang et al., 2024b), overlooking the fine-grained procedural knowledge embedded within the interaction trajectory.
Process Supervision. To address the limitations of outcome-based rewards, particularly the challenge of credit assignment in long-horizon tasks, a growing body of research has explored process supervision (Luo et al., 2024; Xiong et al., 2025). The core idea is to provide agents with more granular feedback at intermediate steps of a task. Early efforts in this area often relied on costly human annotations to label the correctness of each step (Lightman et al., 2023). To automate this, recent methods have proposed various techniques to estimate step-level rewards, such as using Monte Carlo rollouts to predict the future outcome from an intermediate state (Xiong et al., 2024) or training a separate reward model to predict the value of each action (Choudhury, 2025; Wang et al., 2025a). These approaches typically use the estimated step-level rewards to guide the agent via reinforcement learning (Feng et al., 2025; Zhang et al., 2025a) or DPO at the single-action level.
@@ 3 Methodology
3.1 Problem Setting
In this work, HPL follow the same two-stage protocol as previous work (Song et al., 2024; Xiong et al., 2024) that combines a one-shot exploration phase with purely offline preference optimization.
  • Fixed exploration and labeling. A frozen reference policy $\pi_{\text{ref}}$ interacts with the environment once to collect a pool of interaction traces, including both full trajectories and partial segments. We then run Monte Carlo rollouts with $\pi_{\text{ref}}$ on this static pool to derive different level reward estimates, constructing the preference datasets. Importantly, the policy $\pi_{\theta}$ is never updated during this phase, and no further data collection is performed once the datasets are built.
  • Offline preference optimization. Given these fixed datasets, preference-based methods train the target policy $\pi_{\theta}$ using DPO-style objectives at different levels (trajectory, step, and group), without any additional environment interaction or reward queries.
We therefore view our setting as offline preference optimization after a one-shot exploration phase, which is distinct from fully online RL methods such as PPO (Schulman et al., 2017) that continuously interleave policy updates with new environment rollouts throughout training.
@@ 3.3.2 Group-Level Data Generation via Action Group Segmentation
To bridge the gap between coarse trajectories and single actions, we introduce the core concept of an action group. These groups serve as an intermediate unit of reasoning, ideally corresponding to semantically coherent sub-tasks for more effective credit assignment. The generation of group-level data involves two key steps: segmenting trajectories into groups and then estimating a quantitative reward for each group.
@@ 3.3.3 Group-Level Reward Estimation
While step-level MC estimation provides a principled, localized estimate of the expected return at each decision point, in practice the rollout budget per state is small. Under this limited-rollout regime, per-step estimates can have high variance, which makes purely step-level preference learning statistically inefficient. In contrast, our group-level rewards aggregate the outcomes of multiple actions into a single supervision unit, amortizing the same rollout budget over longer sub-trajectories and better capturing their joint contribution to task success.
@@ 3.5 Multi-Granularity Preference Optimization
We briefly analyze the bias-variance properties of this group-level objective in the following proposition, with a full derivation provided in Appendix G.
@@
The final training objective combines these losses:
$\mathcal{L}_{final}^{(s)}=\mathcal{L}_{\text{BC}}+\mathcal{L}_{\text{traj-DPO}}+\mathcal{L}_{\text{step-DPO}}+\mathcal{L}_{\text{group-DPO}}^{(s)},$ (8)
@@ 4.1 Experimental Setup
We evaluate our proposed method, HPL, on three diverse and challenging long-horizon agent benchmarks: ALFWorld (Shridhar et al., 2021), WebShop (Yao et al., 2022), and InterCode-SQL (Yang et al., 2023). In all our benchmarks, the environment returns a single terminal outcome reward at the end of each episode, which defines a finite-horizon episodic MDP and can be modeled with $\gamma$ close to 1 and suitably rescaled rewards. For all experiments, we use Qwen2.5-1.5B-Instruct and Qwen2.5-7B-Instruct as the backbone language models. HPL is compared against a suite of strong baselines, including SFT, RFT (Yuan et al., 2023), ETO (Song et al., 2024), and IPR (Xiong et al., 2024). All methods are initialized from an SFT model trained on expert trajectories generated by a GPT-4o teacher model. For MC estimation, we use $M=5$ rollouts for per group. For Fixed-$N$ and Fixed-$K$ strategy, we set $N=3$ and $K=3$. For Uncertainty strategy, the entropy threshold $\epsilon$ is set to the 80th percentile of all action entropies (i.e., the threshold for the top 20% highest values) computed across the training dataset. Detailed descriptions of the environments, baseline implementations, and all hyperparameters are deferred to Appendix D.
@@ 4.2 Main Results (RQ1, RQ2)
Table 1: Performance comparison of HPL and baselines across agent benchmarks over 3 random seeds. All methods are evaluated using Qwen2.5-1.5B-Instruct and Qwen2.5-7B-Instruct as base models. The best and second-best results are highlighted in bold and with an underline, respectively.
@@
   SFT 60.95±1.09 57.96±1.88 56.56±0.69 26.00±0.50 56.24±0.61 54.33±0.76 52.01±0.43
   RFT (Yuan et al., 2023) 61.19±1.80 60.95±0.86 57.66±1.45 28.17±1.04 58.08±0.64 56.67±0.29 53.79±0.40
   ETO (Song et al., 2024) 65.48±3.60 66.42±2.24 56.57±0.22 28.00±0.87 58.45±1.01 57.67±0.76 55.43±0.86
   IPR (Xiong et al., 2024) 65.24±2.30 66.67±3.68 57.76±1.13 27.83±1.04 58.26±1.78 57.17±1.04 55.49±0.78
@@
   SFT 67.62±2.18 73.63±3.11 60.64±1.12 31.83±1.26 66.70±1.11 65.17±0.76 60.93±0.71
   RFT (Yuan et al., 2023) 71.43±1.89 72.63±3.02 61.16±0.85 33.50±1.00 68.01±0.89 66.33±0.76 62.18±0.11
   ETO (Song et al., 2024) 72.62±2.51 77.86±2.40 61.85±1.00 33.17±1.04 68.32±0.86 67.00±0.50 63.47±0.47
   IPR (Xiong et al., 2024) 73.10±1.80 78.11±3.76 62.01±0.43 33.67±0.58 68.86±1.02 67.17±0.58 63.82±0.69
@@
As shown in Table 1, our HPL framework significantly outperforms all baseline methods across both model scales, providing a strong affirmative answer to our first research question. For the Qwen2.5-7B-Instruct model, our best-performing variant HPL (Semantic) achieves an average score of 67.28, surpassing the strongest single-granularity baselines ETO and IPR by substantial margins of 3.81 and 3.46 points, respectively. This trend of superiority holds true across all three diverse and challenging benchmarks. The advantage of HPL’s hierarchical approach is particularly pronounced in tasks requiring complex, long-horizon generalization. For instance, on the ALFWorld benchmark’s unseen scenarios, HPL (Semantic) achieves a mean success rate of 84.08%, a remarkable improvement of nearly 6 points over the state-of-the-art IPR (78.11%). These results strongly indicate that by integrating preference signals at the trajectory, action, and group levels, HPL effectively resolves the granularity mismatch problem and learns a more robust and generalizable policy.
Our experiments also reveal the critical impact of the action group segmentation strategy, directly addressing our second research question. While all HPL variants consistently outperform the baselines, the adaptive, content-aware segmentation methods generally yield better performance than heuristic approaches. Notably, HPL (Semantic), which leverages a powerful language model to partition trajectories into semantically coherent sub-tasks, consistently emerges as the top-performing variant for both the 1.5B and 7B models. For the 7B model, it outperforms the next-best variant, HPL (Uncertainty), by over 0.35 point in the mean average score. This suggests that the quality of the action groups is paramount; providing the DPO loss with more meaningful, human-aligned sub-tasks as units of comparison leads to a more effective and powerful learning signal. Even so, the strong performance of the simpler Uncertainty and Fixed-N strategies demonstrates the inherent value of incorporating an intermediate granularity, validating the core design of our framework.
@@ 4.3 Analysis of the Curriculum Learning Mechanism (RQ3)
Ablation of Curriculum Components. To quantitatively assess the importance of our dual-layer curriculum, we conduct an ablation study by removing each layer individually, with results presented in Table 2. The primary finding is that the full HPL model, equipped with both curriculum layers, consistently outperforms all ablated variants. Removing the curriculum entirely (HPL Static) results in the most significant performance degradation across both model scales, confirming that employing a curriculum is crucial for effective learning. For the 7B model, HPL Static trails the full HPL by 2.51 average points, a substantial margin that underscores the necessity of a structured learning progression. Furthermore, the results indicate that both the length and difficulty-based curricula contribute positively to the final performance, with their individual removal leading to noticeable performance drops. This demonstrates the synergistic benefit of our dual-layer design, where organizing the learning process first by task complexity (length) and then by solution quality (difficulty) provides a more effective path to mastering complex agent behaviors.